I'm stuck...help.
(x/3) - (x/12) = -4
4x - x = -48
3x = -48
x = -16
As I type it now I think I may be correct. Am I?
Andrea

Answer by longjonsilver(2297)

your method is correct.

To get rid of the fractions in this question we will multiply every term by 12 and then divide the fractions:

(x/3) - (x/12) = -4
12*(x/3) - 12*(x/12) = -4*12
4x - x = -48
3x = -48
x = -48/3
x = -16

CHECK:
(-16/3) - (-16/12)
(-16/3) - (-4/3)
(-16/3) + (4/3)
(-16+4)/3
(-12)/3
-4 --> which is the correct answer, so we KNOW we are correct.

jon.